# Emergent interactions lead to collective frustration in robotic matter

### v1

@@ Abstract
Current artificial intelligence systems show near-human-level capabilities when deployed in isolation. Systems of a few collaborating intelligent agents are being engineered to perform tasks collectively. This raises the question of whether robotic matter, where many learning and intelligent agents interact, shows emergence of collective behaviour. And if so, which kind of phenomena would such systems exhibit? Here, we study a paradigmatic model for robotic matter: a stochastic many-particle system in which each particle is endowed with a deep neural network that predicts its transitions based on the particles’ environments. For a one-dimensional model, we show that robotic matter exhibits complex emergent phenomena, including transitions between long-lived learning regimes, the emergence of particle species, and frustration. We also find a density-dependent phase transition with signatures of criticality. Using active matter theory, we show that this phase transition is a consequence of self-organisation mediated by emergent inter-particle interactions. Our simple model captures key features of more complex forms of robotic systems.
The recent advances in artificial intelligence have brought about systems with near-human-level capabilities in some specific tasks. These advances mostly rely on deep neural networks, which can be trained efficiently using large data sets. In robotics, artificial intelligence is used to give robots the capability to interpret and manipulate their environment [1]. Recently, artificial intelligence systems are being combined to achieve better performance compared to individual systems [2, 3, 4, 5, 6]. In the context of robotics, such collectives contain multiple robotic units that are programmable and interact with their environment. Programmed interactions enable such robot collectives to create different formations [7, 8, 9] and react to global signals to manipulate objects [10].
@@
Robot collectives have been experimentally implemented across spatial scales and system sizes  [11, 12]. Due to the limitations imposed by miniaturization, there is a trade-off between the number of robots in a collective and the capacity of artificial intelligence that can be implemented in individual units. Macroscopic robot collectives comprise, for example, small groups of interacting robots which are each controlled by a performant artificial intelligence [13]. Smaller-scale robots are controlled by microcontrollers and can interact with thousands of robots for self-assembly, collective transport, collective perception, and more [7]. At the micron scale, robots interact via physical or chemical interactions between particles that are either self-driven active particles or particles driven by external forces and controlled via global signals. Such systems have, for example, been implemented as magnetic micro-disks driven by time-varying magnetic fields. Due to their small scale, the capabilities of individual particles in these systems are limited and they are therefore usually controlled externally using global signals.
@@
Conceptually, one can distinguish between different levels on which learning and execution occurs. In decentralised learning and decentralised execution, units learn and behave independently of each other [14]. In the case of centralised learning and decentralised execution, units are controlled by a single machine-learning algorithm, which has been employed in active matter systems to learn chemotactic behaviour [15, 16]. Finally, in the case of centralised learning and decentralised execution, where a single neural network is used by all particles individually but trained using the trajectory of all particles, 2D and 3D active intelligent swimmers moving collectively to reduce their energy expenditure  [17]. Empirical studies in robot collectives have highlighted the emergence of subgroups of robots that may fulfil specialized tasks [18].Such speciation phenomena include the emergence of implicit leaders that provide direction of guidance to other robots. Such experiments have also highlight the emergence of deadlock states, robots perform suboptimal behaviour [19, 20, 21].
@@
Here, we study a paradigmatic theoretical framework for robotic matter that predicts key phenomena of empirical robot collectives. Specifically, we study stochastic many-particle systems in which each particle individually is endowed with artificial intelligence. In these systems the stochastic state transitions are defined by deep neural networks associated with each particle. The predictions by distinct deep neural networks are based on a particle’s environment and they learn by analysing the particle’s trajectory over time. We find that these systems exhibit complex spatio-temporal behaviour: Over time, increasingly complex interactions between particles evolve in a succession of temporal learning regimes. We observe the evolution of multiple particle species and frustration, where the system is locked long-term in a suboptimal state. We also find a density-dependent phase transition with signatures of criticality. Using active matter theory, we show that this phase transition exhibits signatures of criticality and is a consequence of self-organisation mediated by emergent interactions between particles.
@@
Figure 2: Representative trajectories showing the positions of particles over time in distinct temporal regimes (rows). Each column corresponds to a simulation at the specified particle density. For visual purposes we rescaled positions and time in the following way: The x-axis is the lattice position divided by the total lattice size (from left to right 2e4, 2e3, 1e3, 5e2, 4e2, 334 sites, respectively). The y-axis denotes the time in episodes divided by the typical time it takes for a particle to cross the system once. The starting time for each row is indicated in grey and the final time is on the top-right corner of each panel. Line colors show the sign of the average relative velocity of that particle in the last 2e5 episodes of the simulations. Trajectories of representative particles are highlighted in bold. For densities 0.4, 0.5, and 0.6 we have randomly sampled 100 particles to make individual lines distinguishable.
@@ Results
To begin, we consider a stochastic many-particle system where each particle, indexed by $i$, is described by a random variable, $x_{i}$, Fig. 1(a). The state of the system at a given time is then defined by the states of all $N$ particles, $\{x_{i}\}_{i=1,\ldots,N}$. The states of particles are updated according to transition rates between two states $\{x_{i}\}_{i=1,\ldots,N}$ and $\{x_{i}^{\prime}\}_{i=1,\ldots,N}$ per unit time. These transition rates follow from the probabilities of each particle performing individual state transitions $x_{i}\to x_{i}^{\prime}$. In our model, each particle is endowed with a unique deep neural network (DNN) which takes the system configuration, $\{x_{i}\}_{i=1,\ldots,N}$, as an input and computes the single-particle transition rate as an output. Because the system configuration includes the states of other particles, the DNNs may encode effective interactions between particles. DNNs can, in theory, approximate any continuous functions up to any desired accuracy [22], such that the single-particle transition rates may, in principle, encode arbitrarily complex interactions. These interactions may change over time as particles learn from observing their past trajectories and update the parameters of their DNNs based on a reward function. Taken together, the particles in our model are endowed with sensors, processors and actuators  [11]. The model describes a multi-scale system, in which DNNs determine the dynamics of individual particles and, vice versa, collective states on the macroscopic scale influence the learning of the parameters of DNNs. This gives two-way feedback between collective states emerging from effective interactions between particles on the macroscopic scale and the dynamics of DNN parameters,Fig. 1(a).
@@
Here, to investigate the range of possible phenomenologies such systems may exhibit, we study the arguably simplest possible implementation of this framework. In this implementation, the state $x_{i}$ of particles is defined by their position on a one-dimensional lattice of $L$ sites with periodic boundary conditions, Fig. 1(b). A given particle $i$ can hop to the right neighbouring site with probability $p_{i}$ and to the left neighbouring site with probability $1-p_{i}$. This fixes the unit of time. These transition rates are determined by the output layer of feed-forward DNNs which take the lattice environment of size $l$ centered around a given particle as an input. The DNN parameters associated with a given particle are dynamically updated using deep reinforcement learning on the particles’ past trajectory to maximise a discounted reward. The reward function gives a large negative reward when a particle moves to an already-occupied lattice site, and a small positive reward otherwise (Appendix  A).
@@
We used the kinetic Monte Carlo method to perform stochastic simulations of this model  [23]. To reduce the computational cost of training the DNNs, we divided the simulation time into intervals, termed episodes, Fig. 1(c). During each episode, we updated the positions of particles while keeping the DNN parameters fixed. At the end of each episode, we updated the values of the DNN parameters using reinforcement learning. There are different algorithms for training a collection of agents: Multi-agent reinforcement learning (MARL) is used when multiple agents are in a common environment and should learn to cooperate or compete with each other  [24, 25, 26, 27, 28]. Group-agent reinforcement learning (GARL) is used for cases in which each agent operates in a separate environment and only communicates with others to share knowledge [29, 30]. Here, we used an actor-critic reinforcement learning scheme with temporal-difference learning scheme to update the DNN parameters by maximising the cumulative discounted reward calculated from the particle trajectories in each episode  [31] (see Methods). All figures shown here were obtained from simulations, where the number of particles, $N$, was set to 200200200, the local environment size, $l$, was 212121, the system size, $L$, was set depending on particle density, and the discount factor, $\gamma$, was 0.90.90.9. DNNs had a triangular architecture with 21 neurons in the input layers, and 10, 5, and 2 neurons in the successive layers, respectively. We simulated up to $5\cdot 10^{6}$ episodes with learning rate $5\cdot 10^{-4}$. We also tested DNNs with up to five layers, but we observed no qualitative differences in the simulations. To simulate dilute systems without reducing the number of particles, we increased the number of lattice sites to achieve a given particle density while keeping the size of the environment, $l$, fixed. This was to ensure that the DNN sizes and the training dynamics at different particle densities are consistent. This fixes the units of length to the lattice spacing.
@@
From the definition of the model, several questions must be asked that in this form do not arise in active matter systems: First, particles appear in other particles’ environment. Does this lead to the emergence of effective interactions between particles? If so, then, secondly, deep neural networks can, in principle, represent arbitrarily complex interactions that might be qualitatively different for each environment and each particle. Are robotic matter systems amenable to physics-type descriptions? Finally, would effective interactions between particles lead to collective behaviour as in other non-equilibrium systems studied in physics?
@@
Figure 2(a) shows representative trajectories of simulation results for different particle densities and different time windows in a given simulation. Despite the simplicity of the model, we find that particles form complex spatio-temporal structures after a short transient time. These structures change qualitatively over time, indicating that the dynamics transits through different temporal regimes. The spatio-temporal structures also depend on the particle density in that particles at low densities tend to align their velocity in the same direction, while above a density threshold of 0.3 particles at high densities move in opposite directions. In the following, we will analyse these changes as a function of the density more rigorously.
@@
Figure 3: (a) Average velocity of particles relative to the average velocity of all particles in a simulation. Particles are grouped by their asymptotic velocity, which is calculated as an average over the final 2e5 episodes of the simulations. The dashed lines on the time axis denote in this and in all following figure two characteristic times: $t_{1}\approx 1e4$ episodes marks the initial formation of global alignment at low densities; $t_{2}\approx 3e4$ episodes marks the initial formation of two groups of particles moving in opposite direction at high densities. (b) Average absolute velocity of all particles for two representative densities (colours defined in (c)). (c) Log2-fold change of the standard deviation of the number of right-moving particles across simulations. The lower horizontal dashed line is the variance expected if the direction of movement was assigned by chance. The upper dashed line is the maximal variance expected for globally aligned motion. (d) Average reward for two representative densities as a function of time. The inlay is a magnification that shows a decrease in the reward for density 0.5. Colours as in (c).
@@
To this end, we first consider the relative velocity of particles with respect to each other. It is defined by the difference between the individual, instantaneous velocity of a given particle, $v_{i}$, and the average velocity in the system at a given time, $\expectationvalue{v_{i}}$, $v_{i}-\expectationvalue{v_{i}}$. We calculated the relative velocity for each particle and averaged over particles with the same sign of the asymptotic velocity. Figure 3(a) shows that, for low densities, the particles have vanishing relative velocities and therefore are moving in the same direction with a similar absolute velocity. Figure 3(b) shows that this occurs at a characteristic time $t_{1}\approx 1e4$ episods. For densities larger than 0.3, particles form two groups moving in opposite directions. This could reflect a random assignment of the direction of movement of particles. However, the variance of particles moving in a given direction across all simulations is smaller than what is expected from a statistically independent assignment and predicted by the binomial for millions of episodes ($p=0.003$, F-test, Fig. 3(c)). Significantly, this implies that, in the high-density regime, the direction of particles is not chosen randomly, but tuned into a precisely balanced proportion of left- and right-moving particles on the system level.
@@
The fine-tuning of particle directions to a 50-50 split at high densities seems counterintuitive given that they are rewarded for not colliding with other particles. Indeed, while the average reward at low concentrations increases monotonically and reaches an optimal value, at high densities the reward first increases and then decreases at a time that is comparable to the formation of the two groups of particles, Fig. 3(d).
@@
This observation raises two questions: first, what is the reason for the abrupt change in the behaviour of particles? And secondly, why is the high-density regime fine-tuned to a suboptimal reward? A potential answer to these questions could either be that particles learn different behaviours at high and low densities. This should then be reflected in the dynamics of the parameters of the DNNs. An alternative explanation could be that the abrupt change in behaviour emerges from self-organisation due to interactions between particles. In the following, we will aim to understand the origin of this behaviour by investigating in detail the spatio-temporal structures, the time evolution of neural networks, and emergent interactions between particles.
@@
To quantify the time evolution of spatial correlations between particle positions, we calculated the average distance to the nearest-neighbour particle over time, Fig. 4(a). In both density regimes, the nearest-neighbour distance decreases initially with time, showing that particles tend to cluster with other particles. The time evolution of the average nearest-neighbour distance is highly non-monotonic, reflecting different temporal regimes. Despite the formation of clusters, particles nevertheless reduce the probability of overlapping with other particles, Fig. 4(b). To further understand the time evolution of spatial structures we calculated the pair correlation function, which gives the average density of particles as a function of the distance to a given particle, Fig. 4(c). It shows that particles indeed cluster, but the sites adjacent to a given particle tend to be unoccupied. The pair-correlation function also shows that particle-particle correlations are longer in range in the low-density regime compared to the high-density regime.
@@
To understand how the dynamics of the DNN parameters relate to the formation of spatio-temporal structures, we analyse the forces the DNNs encode. Intuitively, interactions between particles are encoded in the temporal change in its velocity, averaged over all environments. To develop this idea, we first represent a given environment by a vector $\vec{e}$, which contains the number of particles at each position in the environment. We now quantify the change in the velocity due to changes in the particle’s environment. To this end, we note that the change in the velocity can be written as the product of changes due to changing environments, and the temporal evolution of the distribution of finding a given environment. Interactions are encoded in the former part, which is given by the directional gradient of the velocity with respect to changes in the environment. In our case, the velocity is determined by the probability of hopping right in a given environment via $2p(\vec{e})-1$. Therefore, the directional derivative of the DNN output for a given input is a natural choice for quantifying interactions. To implement this, for a given environment $\vec{e}_{0}$, we quantified the change in the transition rates $p_{i}$ when the environment $\vec{e}_{0}$ is changed slightly by $\Delta e$. Here, we perturb the environment $\Delta\vec{e}$ by the addition of a single particle. With this, we can quantify interactions by calculating the difference in the neural network output between the original environment and the perturbed environment,
$\frac{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}{\left|\Delta\vec{e}\right|}\,.$ (1)
To quantify the type of interactions more qualitatively, we multiply (1) with the sign of the distance of the particle in $\Delta e$ to obtain a quantity that quantifies the degree with which particles are repulsed by the addition of a new particle to their environment (repulsiveness), $R=-\operatorname{sign}{(\Delta x)}\left[{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}\right]\big{/}{\left|\Delta\vec{e}\right|}$. Negative values of the repulsiveness indicate attractive interactions.
@@
To quantify the so-defined interactions for each particle, we sampled 1000 random environments from the set of environments that host at most 6 particles. We passed these environments, and the corresponding environments that have one additional particle, to the DNNs of each particle at multiple time points, record the DNNs’ output, and calculate the value of $R$, Fig. 6(a). Applying this procedure to all 1000 randomly-chosen environments, particles, and time points over 50 replicates per density leads to data with roughly $10^{7}$dimensions. As a first step towards quantifying interactions between particles, we therefore used a nonlinear dimensionality-reduction technique, uniform manifold approximation, and projection (UMAP), to obtain a first insight into the time evolution of interactions. Particles that are close in these projections tend to behave similarly in interactions. The UMAPs show the systematic time evolution of the response of particles to changes in their environments, Fig. 6(b). This time evolution is qualitatively different in the high and low density regimes, meaning that particles develop different interactions in both regimes, Fig. 6(b). All structures in the UMAPs are independent of the simulation run, Fig. 6(c), showing that particles consistently develop similar possible interactions across runs.
@@
At high densities, we observe the emergence of four clusters of particles in the projected space after a time $2\cdot 10^{5}$ which corresponds to the formation of the two groups of particles moving in different directions, Fig. 2(c). This means that particles evolve into four distinct species, which are distinguished by the way they interact with their environments. Two pairs of these clusters identify particles moving ballistically in a given direction, Fig. 6(c). Particles moving in the same direction further separate into two groups, which interact in different ways depending on higher-order properties of the environments. We were not able to identify simple rules that distinguish between interactions of particles in the clusters corresponding to particles moving in the same direction, showing that particles develop complex interactions that do not follow patterns as in other systems studied in statistical physics.
@@
To gain deeper insights into these interactions, we parametrised environments by the distance of the particle in $e_{j}$ , $\Delta x$, and the distance of the first and second nearest neighbours in $\vec{e}_{0}$, $k_{1}$ and $k_{2}$, respectively. Figure 7(a) shows that interactions are predominantly repulsive with a tendency for their strength to increase over time. This is expected by the choice of the reward function. Additionally, interactions tend be become longer in range over time, as particles obtain non-zero repulsion with respect to other particles at increasing distances. While particles at high densities only interact with the nearest and next-nearest neighbour sites particles at low densities interact over longer distances at large times. Figure 7(a) also shows characteristic branching points, that mark time points after which the response of particles in a given set of environments becomes non-homogeneous. At these time points, particles learn to distinguish an increasing number of different environments.
@@
To investigate how interactions depend on the base environment $\vec{e}_{0}$ we further parametrised the base environment by the distance to the first nearest neighbour, $k_{1}$, and the distance to the second nearest neighbor, $k_{2}$. At low densities, particles first distinguish between base environments, where $k_{1}=1$ and $k_{2}$ is either equal or larger than 2, Figure 7(b). For low densities, after the emergence of nearest-neighbour repulsion, particles first learn to respond to the presence of a particle at a neighbouring site ($\Delta x=1$) depending on the position of the second nearest-neighbour particle. Specifically, particles learn to distinguish between environments, where the second nearest neighbour is two sites away compared to a longer distance. Later, particles learn to respond to the presence of a particle two sites away ($\Delta x=2$) depending on the position of a second nearest neighbour at the same distance. For high densities, after the emergence of nearest-neighbour repulsion, we observe a similar sequence of branching points but at later time points during the simulation.

### v2

@@ Abstract
Current artificial intelligence systems show near-human-level capabilities when deployed in isolation. Systems consisting of collaborating intelligent agents are being engineered to perform tasks collectively. This raises the question of whether robotic matter, where many learning and intelligent agents interact, shows the emergence of collective behaviour. And if so, what kind of phenomena would such systems exhibit? Here, we study a paradigmatic model for robotic matter: a stochastic many-particle system in which each particle is endowed with a deep neural network that optimizes its transitions based on the particles’ environments. For a one-dimensional model, we show that robotic matter exhibits complex phenomena arising from emergent interactions, including transitions between long-lived learning regimes, the emergence of particle species, and frustration. We also find an abrupt, density-dependent change in the behaviour of particles. Using active matter theory, we show that this phenomenon is a reflection of a phase transition with signatures of criticality. Our model captures key phenomena observed in more complex forms of robotic systems and identifies robotic matter as a laboratory for non-equilibrium physics.
Recent advances in artificial intelligence have led to systems with near-human-level capabilities in some specific tasks [1, 2, 3, 4, 5, 6, 7]. These advances rely primarily on deep neural networks, which can be trained efficiently using large datasets. In robotics, artificial intelligence is used to give robots the ability to interpret and manipulate their environment [8]. Recently, artificial intelligence systems have been combined to achieve better performance compared to individual systems [9, 10, 11, 12, 13]. In the context of robotics, such collectives comprise multiple programmable robotic units that interact with their environment. Programmed interactions enable such robot collectives to create different formations [14, 15, 16] and react to global signals to manipulate objects [17].
@@
Robot collectives have been experimentally implemented across spatial scales and system sizes [18, 19, 14, 20, 21]. Due to the limitations imposed by miniaturisation, there is a trade-off between the number of robots in a collective and the capacity of artificial intelligence that can be implemented in individual units. Macroscopic robot collectives comprise, for example, small groups of interacting robots which are each controlled by a performant artificial intelligence [22]. Smaller-scale robots are controlled by microcontrollers and can interact with thousands of robots for self-assembly, collective transport, collective perception, and more [14]. At the micron scale, robots interact via physical or chemical interactions between particles that are either self-driven active particles or particles driven by external forces and controlled via global signals. Such systems have, for example, been implemented as magnetic micro-disks driven by time-varying magnetic fields [23, 24, 25]. Due to their small scale, the capabilities of individual particles in these systems are limited, and they are therefore usually controlled externally using global signals.
@@
Conceptually, one can distinguish between different modes with which learning and execution occur  [26, 27, 28]. In the mode of decentralised learning and decentralised execution, units can learn and behave independently of each other. This mode, for example, has been studied in cooperative and competitive games  [29, 30]. In the mode of centralised learning and decentralised execution, units are controlled by a single machine-learning algorithm. This mode has been used, for example, in active matter systems to study chemotactic behaviour [31, 32, 33, 34], as well as to teach intelligent swimmers to collectively reduce their energy expenditure [35]. Different modes give rise to different phenomenology. Empirical studies in robot collectives have highlighted the emergence of subgroups of robots that may fulfil specialised tasks [36]. Such speciation phenomena include the emergence of implicit leaders that guide other robots. Such experiments have also highlighted the emergence of deadlock states, in which robots perform suboptimal behaviour [37, 38, 39, 40].
@@
Here, we study a paradigmatic theoretical framework for robotic matter that exhibits key phenomena of empirical robot collectives. Specifically, we study stochastic many-particle systems in which each particle has a distinct artificial intelligence. In these systems, the stochastic state transitions are defined by deep neural networks associated with each particle. The optimization done by distinct deep neural networks are based on a particle’s environment, and they learn by analysing the particle’s trajectory over time. Particles, therefore, can learn arbitrarily complex interactions. We find that these systems exhibit complex spatio-temporal behaviour: Over time, increasingly complex interactions between particles evolve in a succession of temporal learning regimes. We observe the evolution of multiple particle species and frustration, where the system is locked long-term in a suboptimal state. Using active matter theory, we show that the system exhibits a density-dependent phase transition, with signatures of criticality, and that this transition is a consequence of self-organisation mediated by emergent interactions between particles.
@@
Figure 2: Representative trajectories showing the positions of particles over time in distinct temporal regimes (rows). Each column corresponds to a simulation at the specified particle density. The y-axis denotes the time in episodes divided by the typical time it takes for a particle to cross the system once. The starting time for each row is indicated in grey, and the final time is in the top-right corner of each panel. Line colours show the sign of the average relative velocity of that particle in the last 2e5 episodes of the simulations. Trajectories of representative particles are highlighted in bold. For densities 0.4, 0.5, and 0.6, we have randomly sampled 100 particles to make individual lines distinguishable.
@@ Results
To begin, we consider a stochastic many-particle system where each particle, indexed by $i$, is described by a random variable, $x_{i}$, Fig. 1(a). The state of the system at a given time is then defined by the states of all $N$ particles, $\{x_{i}\}_{i=1,\ldots,N}$. The states of particles are updated according to transition rates between two states $\{x_{i}\}_{i=1,\ldots,N}$ and $\{x_{i}^{\prime}\}_{i=1,\ldots,N}$ per unit time. These transition rates follow from the probabilities of each particle performing individual state transitions $x_{i}\to x_{i}^{\prime}$. In our model, each particle is endowed with a distinct deep neural network (DNN) which takes the system configuration, $\{x_{i}\}_{i=1,\ldots,N}$, as an input and computes the single-particle transition rate as an output. Because the system configuration includes the states of other particles, the DNNs may encode effective interactions between particles. DNNs can, in theory, approximate any continuous functions up to any desired accuracy [41], such that the single-particle transition rates may, in principle, encode arbitrarily complex interactions. These interactions may change over time as particles learn from observing their past trajectories and update the parameters of their DNNs based on a reward function. Taken together, the particles in our model are endowed with sensors, processors and actuators [18]. The model describes a multi-scale system, in which DNNs determine the dynamics of individual particles and, vice versa, collective states on the macroscopic scale influence the learning of the parameters of DNNs. This gives two-way feedback between collective states emerging from effective interactions between particles on the macroscopic scale and the dynamics of DNN parameters, Fig. 1(a).
@@
Here, to investigate the range of possible phenomenologies such systems may exhibit, we study, arguably, the simplest possible implementation of this framework. In this implementation, the state $x_{i}$ of particles is defined by their position on a one-dimensional lattice of $L$ sites with periodic boundary conditions, Fig. 1(b). A given particle $i$ can hop to the right neighbouring site with probability $p_{i}$ and to the left neighbouring site with probability $1-p_{i}$. This fixes the unit of time. These transition rates are determined by the output layer of feed-forward DNNs, which takes the lattice environment of size $l$ centred around a given particle as an input. The DNN parameters associated with a given particle are dynamically updated using deep reinforcement learning on the particle’s past trajectory to maximise a discounted reward. The reward function gives a large negative reward when a particle moves to an already-occupied lattice site, and a small positive reward otherwise (Appendix A).
@@
We used the kinetic Monte Carlo method to perform stochastic simulations of this model [42]. To reduce the computational cost of training the DNNs, we divided the simulation time into intervals, termed episodes, Fig. 1(c). During each episode, we updated the positions of particles while keeping the DNN parameters fixed. At the end of each episode, we updated the values of the DNN parameters using reinforcement learning. There are different algorithms for training a collection of agents: Multi-agent reinforcement learning (MARL) is used when multiple agents are in a common environment and should learn to cooperate or compete with each other [43, 44, 45, 46, 47]. Group-agent reinforcement learning (GARL) is used for cases in which each agent operates in a separate environment and only communicates with others to share knowledge [48, 49]. Here, we used an actor-critic reinforcement learning scheme with a temporal-difference learning scheme to update the DNN parameters by maximising the discounted reward calculated from the particle trajectories in each episode [28] (see Methods). All figures shown here were obtained from simulations, where the number of particles, $N$, was set to $200$, the local environment size, $l$, was $21$, the system size, $L$, was set depending on particle density, and the discount factor, $\gamma$, was $0.9$. DNNs had a triangular architecture with 21 neurons in the input layer, and $21$, $10$, 55, and $2$ neurons in the successive layers, respectively. We simulated up to $5\cdot 10^{6}$ episodes with learning rate $5\cdot 10^{-4}$. We also tested DNNs with $21$ neurons in each layer, but we observed no qualitative differences in the simulations. To simulate dilute systems without reducing the number of particles, we increased the number of lattice sites to achieve a given particle density while keeping the size of the environment, $l$, fixed. This was done to ensure that the DNN sizes and the training dynamics at different particle densities are consistent. This fixes the units of length to the lattice spacing.
@@
From the definition of the model, several questions arise that in this form do not arise in active matter systems: First, particles appear in other particles’ environment. Does this lead to the emergence of effective interactions between particles? If so, then, secondly, deep neural networks can, in principle, represent arbitrarily complex interactions that might be qualitatively different for each environment and each particle. Are robotic matter systems amenable to physics-type descriptions? Finally, would effective interactions between particles lead to collective behaviour as in other non-equilibrium systems studied in physics?
@@
Figure 2(a) shows representative trajectories of simulation results for different particle densities and different time windows in a given simulation. Despite the simplicity of the model, we find that particles form complex spatio-temporal structures after a short transient time. These structures change qualitatively over time, indicating that the dynamics transition through different temporal regimes. The spatio-temporal structures also depend on the particle density in that particles at low densities tend to align their velocity in the same direction, while above a density threshold of 0.3 particles, at high densities move in opposite directions. In the following, we will analyse these changes as a function of the density more rigorously.
@@
Figure 3: (a) Average velocity of particles relative to the average velocity of all particles in a simulation. Particles are grouped by their asymptotic velocity, which is calculated as an average over the final 2e5 episodes of the simulations. The inlay show the density $0.5$ between episodes $5e3$ and $6e5$. (b) Average absolute velocity of all particles for two representative densities (colours defined in (c)). The dashed lines on the time axis denote in this and in all other figures two characteristic times: $t_{1}\approx 1e4$ episodes mark the initial formation of global alignment at low densities; $t_{2}\approx 3e4$ episodes mark the initial formation of two groups of particles moving in opposite directions at high densities. (c) Log2-fold change of the standard deviation of the number of right-moving particles across simulations. The lower horizontal dashed line is the variance expected if the direction of movement were assigned by chance. The upper dashed line is the maximal variance expected for globally aligned motion. (d) Average reward for two representative densities as a function of time. The inlay is a magnification that shows a decrease in the reward for density 0.5. Colours as in (c).
@@
To this end, we first consider the relative velocity of particles with respect to each other. It is defined by the difference between the individual, instantaneous velocity of a given particle, $v_{i}$, and the average velocity in the system at a given time, $\expectationvalue{v_{i}}$. We calculated the relative velocity for each particle and averaged over particles with the same sign of the asymptotic velocity. Figure 3(a) shows that, for low densities, the particles have vanishing relative velocities and therefore are moving in the same direction with a similar absolute velocity. Figure 3(b) shows that this occurs at a characteristic time $t_{1}\approx 1e4$ episodes. For densities larger than 0.3, particles form two groups moving in opposite directions. This could reflect a random assignment of the direction of movement of particles. However, the variance of particles moving in a given direction across all simulations is smaller than what is expected from a statistically independent assignment of direction ($p=0.003$, F-test, Fig. 3(c)) and lasts for millions of episodes. Significantly, this implies that, in the high-density regime, the direction of particles is not randomly chosen, but tuned into a precisely balanced proportion of left- and right-moving particles on the system scale.
@@
The fine-tuning of particle directions to a balanced 50-50 split at high densities seems counterintuitive, given that they are punished for colliding with other particles. Indeed, while the average reward at low concentrations increases monotonically and reaches almost an optimal value, at high densities the reward first increases and then decreases at a time that is comparable to the formation of the two groups of particles, Fig. 3(d).
@@
This observation raises two questions: first, what is the reason for the abrupt change in the behaviour of particles as the density increases? And secondly, why is the high-density regime fine-tuned to a suboptimal reward? A potential answer to these questions could be that particles learn different behaviours at high and low densities. This should then be reflected in the dynamics of the parameters of the DNNs. An alternative explanation could be that the abrupt change in behaviour emerges from self-organisation due to interactions between particles. In the following, we will aim to understand the origin of this behaviour by investigating in detail the spatio-temporal structures, the time evolution of neural networks, and emergent interactions between particles.
@@
To quantify the time evolution of spatial correlations between particle positions, we calculated the average distance to the nearest-neighbour particle over time, Fig. 4(a). In both density regimes, the nearest-neighbour distance decreases initially with time, showing that particles tend to cluster with other particles. However, the time evolution of the average nearest-neighbour distance is highly non-monotonic, reflecting different temporal regimes. Despite the formation of clusters, particles nevertheless reduce the probability of overlapping with other particles, Fig. 4(b). To further understand the time evolution of spatial structures, we calculated the pair correlation function, which gives the average density of particles as a function of the distance to a given particle, Fig. 4(c). It shows that particles indeed cluster, but the sites adjacent to a given particle tend to be unoccupied. The pair-correlation function also shows that particle-particle correlations are longer in range in the low-density regime compared to the high-density regime.
@@
To understand interactions between particles, for a given environment $\vec{e}_{0}$, we quantified the change in the transition probabilities $p_{i}$ when the environment $\vec{e}_{0}$ changes slightly by $\Delta e$. Here, we perturb the environment $\vec{e}_{0}$ by adding a single particle. With this, we can quantify interactions by calculating the difference in the neural network output between the original environment and the perturbed environment (see Appendix G),
$\nabla_{\Delta\vec{e}}p(\vec{e_{0}})\approx\frac{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}{\left|\Delta\vec{e}\right|}.$ (1)
To quantify the type of interactions more qualitatively, we multiply (1) with the sign of the distance of the particle in $\Delta e$ to obtain a quantity that quantifies the degree with which particles are repulsed by the addition of a new particle to their environment, which we term repulsiveness, $R=-\operatorname{sign}{(\Delta x)}\left[{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}\right]\big/{\left|\Delta\vec{e}\right|}$. Negative values of the repulsiveness indicate attractive interactions.
@@
To quantify the so-defined interactions for each particle, we sampled 1000 random environments from the set of environments that host at most 6 particles. We passed these environments, and the corresponding environments that have one additional particle, to the DNNs of each particle at multiple time points, record the DNNs’ output, and calculate the value of $R$, Fig. 6(a). Applying this procedure to all 1000 randomly-chosen environments, particles, and time points over 50 replicates per density leads to data with roughly $10^{7}$dimensions. As a first step towards quantifying interactions between particles, we therefore used a nonlinear dimensionality-reduction technique, uniform manifold approximation, and projection (UMAP), to obtain a first insight into the time evolution of interactions. Particles that cluster these projections tend to behave similarly in interactions. The UMAPs show the systematic time evolution of the response of particles to changes in their environments, Fig. 6(b). This time evolution is qualitatively different in the high and low-density regimes, meaning that particles develop different interactions in both regimes, Fig. 6(b). All structures in the UMAPs are independent of the simulation run, Fig. 6(c), showing that particles consistently develop similar possible interactions across runs.
@@
At high densities, we observe the emergence of four clusters of particles in the projected space after a time $2\cdot 10^{5}$, which corresponds to the formation of the two groups of particles moving in different directions, as shown in Fig. 2(c). This means that particles evolve into four distinct species that are distinguished by the way they interact with their environments. Two pairs of these clusters identify particles that move ballistically in a given direction, Fig. 6(c). Particles moving in the same direction further separate into two groups, which interact in different ways depending on higher-order properties of the environments. We were unable to identify simple rules that distinguish between interactions of particles in the clusters corresponding to particles moving in the same direction, showing that particles develop complex interactions that do not follow patterns as in other systems studied in statistical physics. The emergence of groups of particles that interact differently resembles speciation phenomena in evolutionary biology.
@@
To gain deeper insights into these interactions, we parametrised environments by the distance of the particle in $e_{j}$, $\Delta x$, and the distance of the first and second nearest neighbours in $\vec{e}_{0}$, $k_{1}$ and $k_{2}$, respectively. Figure 7(a) shows that interactions are predominantly repulsive with a tendency for their strength to increase over time. This is expected by the choice of the reward function. Additionally, interactions tend to become longer in range over time, as particles obtain non-zero repulsion with respect to other particles at increasing distances. While particles at high densities only interact with the nearest and next-nearest neighbour sites, particles at low densities interact over longer distances at large times. Figure 7(a) also shows characteristic branching points that mark time points after which the response of particles in a given set of environments becomes non-homogeneous. At these time points, particles learn to distinguish an increasing number of different environments.
@@
To investigate how interactions depend on the base environment $\vec{e}_{0}$ we further parametrised the base environment by the distance to the first nearest neighbour, $k_{1}$, and the distance to the second nearest neighbour, $k_{2}$. At low densities, particles first distinguish between base environments, where $k_{1}=1$ and $k_{2}$ is either equal to or larger than 2, Fig. 7(b). For low densities, after the emergence of nearest-neighbour repulsion, particles first learn to respond to the presence of a particle at a neighbouring site ($\Delta x=1$) depending on the position of the second nearest-neighbour particle. Specifically, particles learn to distinguish between environments, where the second nearest neighbour is two sites away compared to a longer distance. Later, particles learn to respond to the presence of a particle two sites away ($\Delta x=2$), depending on the position of a second nearest neighbour at the same distance. For high densities, after the emergence of nearest-neighbour repulsion, we observe a similar sequence of branching points but at later time points during the simulation.